OpenStudy (anonymous):

How many turns needed for a rider to travel one mile on a Ferris wheel that is 250 feet in diameter and height of 264?

OpenStudy (phi):

If you knew that if you go one turn on the wheel you travel 785.4 ft, can you figure out how many turns to go 5280 ft? Use a ratio: \[ \frac{1 \text{ turn}}{785.4 \text{ ft}} = \frac{x \text{ turns}}{5280\text{ ft}}\] btw, you figure out the distance of one turn by using \[ C= 2\pi r = \pi d \]